# Accurate Estimation of Mutual Information in High Dimensional Data

### v1

@@ Abstract
Mutual information (MI) is a measure of statistical dependencies between two variables, widely used in data analysis. Thus, accurate methods for estimating MI from empirical data are crucial. Such estimation is a hard problem, and there are provably no estimators that are universally good for finite datasets. Common estimators struggle with high-dimensional data, which is a staple of modern experiments. Recently, promising machine learning-based MI estimation methods have emerged. Yet it remains unclear if and when they produce accurate results, depending on dataset sizes, statistical structure of the data, and hyperparameters of the estimators, such as the embedding dimensionality or the duration of training. There are also no accepted tests to signal when the estimators are inaccurate. Here, we systematically explore these gaps. We propose and validate a protocol for MI estimation that includes explicit checks ensuring reliability and statistical consistency. Contrary to accepted wisdom, we demonstrate that reliable MI estimation is achievable even with severely undersampled, high-dimensional datasets, provided these data admit accurate low-dimensional representations. These findings broaden the potential use of machine learning-based MI estimation methods in real-world data analysis and provide new insights into when and why modern high-dimensional, self-supervised algorithms perform effectively.
aDepartment of Physics  bDepartment of Biology  cInitiative in Theory and Modeling of Living Systems
Emory University, Atlanta, GA 30322, USA
@@ 1 Introduction
Mutual information (MI) is a fundamental measure of statistical dependence between two variables [1]. It captures both linear and nonlinear associations, is invariant under reparameterizations, and is zero if and only if the variables are statistically independent. These properties make MI a key tool across disciplines, from neuroscience to computer vision [2]. In systems neuroscience, MI estimation plays an important role in understanding neural coding and decoding processes, analyzing spike trains in single neurons and neural populations, and studying patterns of information transfer across brain areas and behaviors [3, 4, 5, 6, 7, 8, 9]. Similarly, in brain imaging studies, MI quantifies functional connectivity between brain regions, illuminating cognitive processes and effects of neurological disorders [10, 11, 12]. MI has also proven useful in domains like protein sequence alignment and contact prediction [13, 14, 15], and in the inference of gene regulatory networks [16]. In computer vision, state-of-the-art models, such as CLIP [17], utilize loss functions related to MI to align visual and textual representations. Similarly, self-supervised learning frameworks such as Barlow Twins [18] can be interpreted as mutual information-based objectives designed to enforce cross-modal correspondences [19].
[…]
@@
Recent advances in machine learning have introduced neural network (NN)-based estimators for MI, which aim to circumvent the limitations of traditional methods. These estimators frame MI estimation as optimization over a family of functions [32, 33, 34, 35, 36, 37]. In principle, they can work even for very high-dimensional data, producing estimates in scenarios that are out of reach for traditional methods. For instance, they supposedly can compute MI between images, with the dimensionality of thousands. However, the practical accuracy of these methods remains unclear. First, most of them have been tested primarily on synthetic data with simple dependence structures and essentially infinite sample sizes, raising questions about their real-world applicability. Second, since universally good MI estimation is impossible, internal consistency checks are essential to signal whether the output can be trusted. Such checks are not widely adopted. Third, NN estimators depend on hyperparameters, such as criteria for stopping training, and how to choose these parameters remains unclear.
@@
Here, we systematically address these gaps. We argue that successful neural MI estimation requires: (i) even high-dimensional data having a low-dimensional latent structure; (ii) the critic sufficiently expressive (e.g., embedding dimension and nonlinearity) to capture it; and (iii) sufficient data to resolve statistical dependencies in the latent (not full data) space. Our specific contributions include: (1) development of max-test heuristic for early stopping to mitigate overfitting in MI estimation; (2) introduction of probabilistic critics, which improve estimation for large MI values; (3) unification of correlation-based and neural MI estimators under the Donsker-Varadhan [38] framework; (4) benchmarking estimators on synthetic and real-world data, demonstrating when and why MI estimation succeeds; and (5) development of a practical guide for MI network estimation.
@@ 2 Background
Traditional MI Estimation Methods.
Estimating mutual information (MI) from finite data is notoriously challenging, especially for high-dimensional continuous variables [39]. To illustrate this, consider a simple argument: suppose each component of $X$ and $Y$ lies within a bounded range of size $A$, and the joint density $p(x,y)$ is smooth, with its smallest feature on the scale of $a$. Then accurate MI estimation requires $N\gg(A/a)^{K_{\text{tot}}}$ samples, where $K_{\text{tot}}=K_{X}+K_{Y}$ is the total dimensionality of $X$ and $Y$. Even when $A/a$ is only slightly greater than 1, the required sample size is exponential in $K_{\text{tot}}$, illustrating the classic curse of dimensionality. The situation is even worse when the coordinate system in which $p$ is smooth is unknown, or when the variables are unbounded. Thus, while many methods have been developed to estimate MI from continuous data, they often break down beyond $K_{\rm tot}\sim 10$ dimensions [40] (cf. SI Fig. 10). In contrast, most modern datasets are high-dimensional, e.g., images with thousands of pixels, or neural recordings from thousands of units.
@@
The inability of traditional methods to overcome the curse of dimensionality has motivated neural network (NN)-based approaches. Deep NNs can capture complex nonlinear dependencies in high-dimensional data and, in some cases, generalize well even from modest sample sizes [41]. Neural variational methods have become particularly influential for MI estimation [32, 33, 34, 38]. Specifically, typically, we do not have access to the full joint distribution $p(x,y)$ or the marginals $p(x)$ and $p(y)$, but we can draw samples from them. Variational estimators leverage this by reformulating MI in terms of a Kullback-Leibler (KL) divergence:
@@
Then using, the Donsker–Varadhan (DV) representation of the first KL divergence in Eq. (1) [38], $D_{\text{KL}}(P\parallel Q)\geq\max_{T}\left\{\mathbb{E}_{P}[T]-\log\mathbb{E}_{Q}[e^{T}]\right\}$, learning the critic function $T(x,y)$ via a NN, and replacing expectations with sample averages, one obtains the MINE estimator [42, 35] (Appx. Eq. 7). However, MINE suffers from high variance, and its estimate is not a strict lower bound on MI when the normalization term is approximated by Monte Carlo sampling [35]. One proposed way to address these problems is to clip the critic to the range ±$\pm\tau$, yielding the SMILE estimator [36] (Appx. Eq. 8). Alternatively, applying the DV representation to the second KL divergence in Eq. (1) instead leads to the InfoNCE estimator for MI [37], widely used in contrastive learning (Appx. Eq. 11). See derivations for all the methods in Appx. A.1.1.
Although early studies employed simple critic networks, the architecture and expressivity of $T(x,y)$ strongly influence estimator performance. We return to this in Sec. 3.
Limitations of Existing Neural Estimators.
Despite their popularity, it is unclear if existing neural MI estimators are truly accurate, calling their widespread use into question. First, most quantitative tests of the estimators to date have been performed on synthetic data with low dimensionality (e.g., $K_{X},K_{Y}\sim 10$), where the true MI is known analytically. However, traditional estimators such as kNN-based methods [21, 40] already perform well in this regime, even for non-Gaussian data [31]. Unless neural estimators clearly outperform these simpler methods in higher-dimensional settings (e.g., $K\gtrsim 100$), their practical value is limited. Yet, evaluations in this regime remain scarce.
Second, when $X$ and $Y$ are jointly Gaussian, MI can be computed exactly from their correlation matrix. Since correlation matrices can be reliably estimated when $K/N\ll 1$ [43, 44], this provides a natural benchmark. If a neural estimator fails when a linear method succeeds, it is evidently not exploiting all the statistical structure in the data. Nevertheless, such comparisons are rarely reported. As we will demonstrate below, some neural estimations fall short on this metric.
Finally, many studies validate estimators in effectively infinite-sample regimes [35, 36], where a fresh data batch is drawn at every training step. This sidesteps overfitting and gives an overly optimistic view of estimator performance. In real applications, sample size is often small ($N\sim K$), and success for infinite-data does not imply practical utility. We will show that unbiased estimation is sometimes possible even in this heavily undersampled regime, but only if data have a simple latent structure.
@@ 3 A Generalized Critic
Joint (Concatenated) Critic: Setting $g$ and $h$ to the identity maps and letting $f$ be a neural network that operates on the concatenated inputs reproduces the joint-critic architecture used in MINE [42].
(Deep) Separable Critic: Choosing $f(g,h)=\langle g,h\rangle$, with $g$ and $h$ implemented as vector-valued embeddings (e.g., via multilayer perceptrons), yields the separable critic used in InfoNCE [37].
Although certain critics are commonly paired with specific objectives, this pairing is not obligatory. For instance, separable critics, typically used with contrastive losses, can be combined with non-contrastive objectives. Likewise, joint critics can be used in contrastive settings.
@@
Concatenated Quadratic Critic: We show that choosing $g$ and $h$ as linear projections (e.g., identity mapping) and $f$ as a quadratic form of its concatenated arguments reduces MI estimation, denoted here as $I_{\rm CCA}(X;Y)$, to measuring canonical correlations [45] in a shared low-dimensional space. See Appx. A.1.2 for details.
Probabilistic Critic (VSIB Framework): Rather than using deterministic mappings, $f$, $g$, and $h$ can be stochastic, leading to variational objectives that regularize embedding distributions. We implement this with a loss similar to that introduced in [19] (see Appx. A.2 for details):
@@ 4.1 Estimator Performance in the Infinite Data Regime
We first evaluate neural MI estimators in an idealized, effectively infinite-sample setting that eliminates overfitting. Although common in prior work, this regime obscures many challenges encountered with finite, real-world data. To benchmark performance, we use synthetic datasets of both low and high dimensionality with known ground-truth mutual information. Throughout the paper, we adopt the following notation. $X$ and $Y$ are the observed variables whose MI we estimate; $Z_{X}$ and $Z_{Y}$ denote their low-dimensional embeddings by a critic. The observed dimensionalities are $K_{X}$ and $K_{Y}$ ($K$ if both are equal), while $K_{Z}$ is the true latent dimensionality when it differs from $K$. Finally, $k_{Z}$ is the embedding dimensionality of the critic, which may differ from $K_{Z}$.
@@
Figure 1: MI estimators in the low-dimensional, infinite-data regime. Each panel plots running MI estimates over training iterations for five true MI levels (increasing every 4000 iterations). Each step introduces a fresh batch of 128 samples. We compare the CCA-based estimator (optimal for Gaussian data), InfoNCE, SMILE, and their probabilistic variants (denoted with VSIB). Faint curves show raw estimates; bold curves show smoothed trends (see Appx. A.6). Left: For jointly Gaussian $X,Y$, all estimators initially perform well. InfoNCE plateaus at its well-known intrinsic upper bound [37] $\log$(batch size) $\approx 7$ bits, while SMILE begins to overestimate at high MI, indicating overfitting. $I_{\rm CCA}$ overlaps with ground truth, as expected. Middle: Cubing $Y$ breaks linearity of correlations, and $I_{\rm CCA}$ fails. Nonetheless, InfoNCE behavior is almost unchanged, and SMILE remains reasonably effective with sufficient training (see SI Fig. 9). Both slightly underestimate at low MI, and for SMILE this is largely offset by its intrinsic positive bias at high MI. Right: Passing $X$ and $Y$ through separate frozen teacher networks (one hidden softplus layer, 1024 units) creates highly nonlinear dependencies. All estimators underestimate MI.
@@
We begin with low-dimensional data ($K_{X}=K_{Y}=10$), where both traditional (e.g., kNN) and neural estimators are expected to perform well. We first analyze jointly Gaussian data with a chosen correlation. It is known since 1950s that, in this case, MI estimate reduces to the sum of information in each canonical correlation pair [46, 47, 48] $I_{\rm CCA}=-\frac{1}{2}\sum_{i}^{K_{Z}}\log(1-\rho_{i}^{2})$, where $\rho_{i}$ is the canonical correlation. We then increase complexity by (i) cubing $Y$ or (ii) passing both variables through separate fixed teacher networks, thereby introducing diverse nonlinear dependencies.
@@
Figure 1 (left) confirms that InfoNCE, SMILE, and CCA track the true MI on Gaussian data (until InfoNCE saturates [37]). The data matrices used here are aggregated across all batches seen during training at a given MI level. If the critic is expressive enough, SMILE is expected to show high variance and to overestimate at high MI since it evaluates log-sum-exp in its DV bound, Eq. (7), which is biased due to the Jensen’s inequality [49, 50]. Indeed, we observe both problems, but it is encouraging that neither is strong, at least at low MI.
@@
Cubing $Y$ introduces nonlinear correlations, so that $I_{\rm CCA}$ fails, cf. Fig. 1 (middle). In contrast, neural estimators recover MI after sufficient training (see SI Fig. 9). Both exhibit a small but significant negative bias at low MI, likely due to the reduced invertibility of the cubic map near $y=0$. When $X$ and $Y$ are passed through separate frozen teacher networks, Fig. 1 (right), all estimators underestimate MI, although neural methods still outperform the CCA baseline. We again attribute this bias to reduced invertibility, now due to the softplus saturation in the teacher networks.
@@
SI Fig. 8 surveys other neural estimators and confirms that InfoNCE and SMILE consistently outperform the rest. Therefore, in what follows, we restrict our analysis to these two methods, using a fixed clipping factor $\tau=5$ for SMILE. Combined, these results validate MI neural estimators—specifically InfoNCE, SMILE, and their VSIB variants—in low dimensions with abundant data. However, they still underestimate MI for strongly nonlinear dependencies, even with effectively unlimited data. Thus, seemingly, there is little reason to prefer them over simpler correlation-based or kNN methods, which already perform well in these scenarios even for smaller datasets [21, 40, 31].
@@
Figure 2: MI estimators in the high-dimensional, infinite-data regime. We extend the setup of Fig. 1 to $K_{X}=K_{Y}=500$, embedding $K_{Z}=10$ latent variables into high-dimensional $X$ and $Y$. Left: A linear transformation (e.g., identity and replication) expands $Z$ to 500-dimensional $X$ and $Y$. $I_{\rm CCA}$ with $k_{z}\geq K_{Z}=10$ dimensions still accurately recovers the ground truth. Middle: A frozen nonlinear teacher network maps $Z$ to 500-dimensional $X$ and $Y$. Unlike the linear case, $I_{\rm CCA}$ fails due to the nonlinearity of the transformation. Increasing $k_{Z}>K_{Z}$ detects spurious correlations, inflating MI estimates and illustrating the limitations of linear methods in nonlinear settings. Right: Neural estimators (InfoNCE, SMILE, and their VSIB variants) are applied directly to the full 500-dimensional data. All are accurate across the full range of true MI values, performing even better than in Fig. 1 due to improved invertibility of the nonlinear transformation in high dimensions.
High-Dimensional $X$ and $Y$.
@@
To distinguish effects of the observed vs. latent dimensionality, we increase the former, while keeping the latter fixed. For this, we start again with the jointly Gaussian 10-d $X$ and $Y$ with the known ground truth MI. We then (i) replicate each of the ten components of $X$ and $Y$ 50 times (denoted as $X\to{\mathbf{I}}^{(50)\otimes}[X]$), and (ii) pass $X$ and $Y$ through distinct frozen teacher networks, embedding each of them into 500 dimensions (denoted as $X\to F_{X}^{(500)}[X]$, and similarly for $Y$). Both cases result in $K_{X}=K_{Y}\equiv K=500$, while $K_{Z}=10$. The former case has only linear correlations, while the latter has many nonlinear ones.
@@
Figure 2 shows that, in the linear replication setup, $I_{\rm CCA}$ accurately recovers the true MI when $k_{Z}\geq K_{Z}$. In the nonlinear case, the CCA-based approach breaks down: as the number of detected canonical pairs $k_{Z}$ increases, the method approximates nonlinear dependencies with an ever larger set of linear projections, inflating the MI estimate with no obvious upper bound. In contrast, neural estimators (InfoNCE, SMILE, and their VSIB variants) recover the ground truth MI when applied directly to the full $K=500$-dimensional spaces (up to InfoNCE’s saturation). Performance even surpasses the low-dimensional case, Fig. 1, because non-invertible, saturated softplus regions in one random embedding can be inverted in others, allowing reconstruction of the full latent manifold.
@@
These results highlight two points. First, neural estimators may work well in high-dimensional nonlinear settings, where traditional estimators break down [40] (See SI Fig. 10). Second, matching the estimator to the statistical structure of the data (e.g., using $I_{\rm CCA}$ for high-dimensional linear correlations) may result in statistically and computationally more efficient estimation (e.g., linear approaches can recover MI with far fewer samples [51]).

### v2

@@ Abstract
Mutual information (MI) is a fundamental measure of statistical dependence between two variables, yet accurate estimation from finite data remains notoriously difficult. No estimator is universally reliable, and common approaches fail in the high-dimensional, undersampled regimes typical of modern experiments. Recent machine learning–based estimators show promise, but their accuracy depends sensitively on dataset size, structure, and hyperparameters, with no accepted tests to detect failures. We close these gaps through a systematic evaluation of classical and neural MI estimators across standard benchmarks and new synthetic datasets tailored to challenging high-dimensional, undersampled regimes. We contribute: (i) a practical protocol for reliable MI estimation with explicit checks for statistical consistency; (ii) confidence intervals (error bars around estimates) that existing neural MI estimator do not provide; and (iii) a new class of probabilistic critics designed for high-dimensional, high-information settings. We demonstrate the effectiveness of our protocol with computational experiments, showing that it consistently matches or surpasses existing methods while uniquely quantifying its own reliability. We show that reliable MI estimation is sometimes achievable even in severely undersampled, high-dimensional datasets, provided they admit accurate low-dimensional representations. This broadens the scope of applicability of neural MI estimators and clarifies when such estimators can be trusted.
@@ 1 Introduction
Mutual information (MI) is a fundamental measure of statistical dependence between two variables (Shannon, 1948). It captures both linear and nonlinear associations, is invariant under reparameterizations, and is zero if and only if the variables are statistically independent. This makes MI a key tool across disciplines, from neuroscience to computer vision (Viola & Wells III, 1997). In systems neuroscience, MI estimation plays an important role in understanding neural coding, analyzing spike trains in single neurons and neural populations, and studying patterns of information transfer across brain areas and behaviors (Tang et al., 2014; Palmer et al., 2015; Panzeri et al., 2001; Strong et al., 1998; Pica et al., 2017; Runyan et al., 2017; Pascual et al., 2024). Similarly, in brain imaging, MI quantifies functional connectivity between brain regions, illuminating effects of neurological disorders (Ince et al., 2017; Hlinka et al., 2011; Li, 2022). MI has also proven useful in domains like protein sequence alignment and contact prediction (Marks et al., 2011; Lu et al., 2020; Sgarbossa et al., 2023), and in the inference of gene regulatory networks Margolin et al. (2006). In computer vision, state-of-the-art models, such as CLIP (Radford et al., 2021), utilize loss functions related to MI to align visual and textual representations. Similarly, self-supervised learning frameworks such as Barlow Twins (Zbontar et al., 2021) can be interpreted as mutual information-based objectives designed to enforce cross-modal correspondences (Abdelaleem et al., 2025).
For continuous variables $X$ and $Y$, MI is $I(X;Y)=\int dx\ dy\,p(x,y)\log_{2}\frac{p(x,y)}{p(x)p(y)}$ (measured in bits [^1] Most estimators are naturally expressed in nats, where $\log$ is the natural logarithm. Thus in all derivations in this paper $\log=\ln$. However, when reporting the MI values, we convert to bits for clarity, $\log=\log_{2}$.), where $x$ and $y$ are specific values of the variables, $p(\cdot)$ are the corresponding probability densities, and the integration is over the domain of the variables. (For discrete $X$ and $Y$, sums over distributions are used instead.) Since MI is a nonlinear function of $p(x,y)$, substituting an unbiased estimate of $p$ into the definition of MI results in a biased estimate of MI. Typically, the bias of estimators is a more serious problem than their variance, particularly for continuous variables, because MI is invariant under reparameterization, while it is impossible to construct an estimator that is covariant under all reparameterizations (Holy & Nemenman, 2002). Traditional methods, such as histogram-based, k-nearest neighbors (kNN), box-counting, or kernel-based (Kraskov et al., 2004; Gao et al., 2015; Ross, 2014; Amir Haeri & Ebadzadeh, 2014; Khan et al., 2007; Steuer et al., 2002; Daub et al., 2004; Trappenberg et al., 2005; Fransens et al., 2004), struggle to reduce the bias for high-dimensional data since they require the number of samples that grows exponentially with the dimensionality (Walters-Williams & Li, 2009; Khan et al., 2007; Gao et al., 2015; Czyz et al., 2023).
@@
Recent advances in machine learning resulted in neural network (NN)-based estimators for MI, which aim to circumvent the limitations of traditional methods. These estimators frame MI estimation as optimization over a family of functions (Barber & Agakov, 2003; 2004; Nguyen et al., 2010; Poole et al., 2019; Song & Ermon, 2019; van den Oord et al., 2018). In principle, they can work even for very high-dimensional data that are out of reach for traditional methods. For instance, they supposedly compute MI between images, with the dimensionality of thousands. However, the practical accuracy of these methods remains unclear. First, most of them have been tested primarily on synthetic data with simple dependence structures and unrealistically large datasets. Second, since universally good MI estimation without smoothness assumptions on the underlying distribution is impossible (see, e.g., Paninski (2003); Kandasamy et al. (2015)), internal consistency checks are essential to signal whether the output can be trusted. Such checks are not widely adopted. Third, NN estimators depend on hyperparameters, such as criteria for stopping training, and optimal parameter choices are unclear.
@@
Here, we systematically address these gaps. We argue that successful neural MI estimation requires: (i) even high-dimensional data having a low-dimensional latent structure; (ii) the critic sufficiently expressive (e.g., embedding dimension and nonlinearity) to capture it; and (iii) sufficient data to resolve statistical dependencies in the latent (not full data) space. Our specific contributions include: (1) developing a practical protocol for MI estimation, which includes an early-stopping heuristic to minimize estimation bias, internal bias checks, an estimation of error bars (confidence intervals)—the latter is particularly rare among MI estimators; (2) introducing probabilistic critics, which improve estimation for large MI values; (3) benchmarking our approach against other estimators on synthetic and real-world data, demonstrating when and why MI estimation succeeds.
@@ 2 Background
Traditional MI Estimation.
Estimating mutual information (MI) from finite data is notoriously challenging, especially for high-dimensional continuous variables (Paninski, 2003). To see this, consider a simple argument: suppose each component of $X$ and $Y$ lies within a bounded range of size $A$, and the joint density $p(x,y)$ is smooth, with its smallest feature on the scale of $a$. Then accurate estimation requires $N\gg(A/a)^{K_{\text{tot}}}$ samples, where $K_{\text{tot}}=K_{X}+K_{Y}$ is the total dimensionality of $X$ and $Y$. If $A/a>1$, the required sample size is exponential in $K_{\text{tot}}$, illustrating the classic curse of dimensionality. The situation is even worse when the coordinate system in which $p$ is smooth is unknown (Holy & Nemenman, 2002), or when the variables are unbounded. Thus, while many methods have been developed to estimate MI, they often break down beyond $K_{\rm tot}\sim 10$ dimensions Holmes & Nemenman (2019) (cf. Appx. Fig. 9). In contrast, most modern datasets are high-dimensional, e.g., images with thousands of pixels, or neural recordings from thousands of units.
@@
The struggle against the curse of dimensionality has motivated neural network (NN)-based approaches since deep NNs can capture complex nonlinear dependencies in high-dimensional data (LeCun et al., 2015). Neural variational methods have become particularly influential for MI estimation (Barber & Agakov, 2003; 2004; Nguyen et al., 2010; Donsker & Varadhan, 1983). Specifically, typically, we do not have access to the full joint distribution $p(x,y)$ or the marginals $p(x)$ and $p(y)$, but we can draw samples from them. Variational estimators leverage this by reformulating MI in terms of a Kullback-Leibler (KL) divergence:
@@
Then using, the Donsker–Varadhan (DV) representation of the first KL divergence in Eq. (1) (Donsker & Varadhan, 1983), $D_{\text{KL}}(P\parallel Q)\geq\max_{T}\left\{\mathbb{E}_{P}[T]-\log\mathbb{E}_{Q}[e^{T}]\right\}$, learning the critic function $T(x,y)$ via a NN, and replacing expectations with sample averages, one obtains the MINE estimator (Belghazi et al., 2018; Poole et al., 2019) (Appx. Eq. 7). However, MINE suffers from high variance, and its estimate is not a strict lower bound on MI when the normalization term is approximated by Monte Carlo sampling (Poole et al., 2019). One addresses this by clipping the critic to the range $\pm\tau$, yielding the SMILE estimator (Song & Ermon, 2019) (Appx. Eq. 8). Alternatively, applying the DV representation to the second KL divergence in Eq. (1) instead leads to the InfoNCE estimator (van den Oord et al., 2018), widely used in contrastive learning (Appx. Eq. 11). See derivations for all the methods in Appx. A.1.1. Although early studies employed simple critic networks, the architecture and expressivity of $T(x,y)$ strongly influence estimator performance. We return to this in Sec. 3.
Limitations of Neural Estimators.
Despite their popularity, it is unclear if existing neural MI estimators are truly accurate, calling their widespread use into question. First, most tests (Czyz et al., 2023) of the estimators to date have been performed on synthetic data with low dimensionality (e.g., $K_{X},K_{Y}\sim 10$), where the true MI is known analytically. However, traditional estimators such as kNN-based methods (Kraskov et al., 2004; Holmes & Nemenman, 2019) already perform well then (Czyz et al., 2023). Unless neural estimators clearly outperform these simpler methods in high dimensions $K\gtrsim 100$), their practical value is limited. Yet, evaluations in this regime remain scarce.
Second, when $X$ and $Y$ are jointly Gaussian, MI can be computed exactly from their correlation matrix. Since correlation matrices can be reliably estimated when $K/N\ll 1$ (Bouchaud et al., 2007; Swain et al., 2025), this provides a natural benchmark. If a neural estimator fails when a linear method succeeds, it is not exploiting all the statistical structure in the data. Nevertheless, such comparisons are rare. As we will demonstrate below, some neural estimations fall short on this metric.
Finally, estimators are often validated in effectively infinite-sample regimes (Poole et al., 2019; Song & Ermon, 2019), with a fresh data batch at every training step. This sidesteps overfitting and gives an overly optimistic view of estimator performance. In real life, sample size is often small ($N\sim K$), and success for infinite-data does not imply practical utility. We will show that unbiased estimation is sometimes possible even in this heavily undersampled regime if data have a simple latent structure.
@@ 3 A Generalized Critic
Joint (Concatenated) Critic: Setting $g$, $h$ to the identity maps and letting $f$ be a NN that operates on the concatenated inputs reproduces the joint-critic architecture used in MINE (Belghazi et al., 2018).
(Deep) Separable Critic: Choosing $f(g,h)=\langle g,h\rangle$, with $g$, $h$ being vector-valued embeddings (e.g., via multilayer perceptrons), yields the separable critic of InfoNCE (van den Oord et al., 2018).
Although certain critics are commonly paired with specific objectives, this pairing is not obligatory. For instance, separable critics, typically used with contrastive losses, can be combined with non‑contrastive objectives. Likewise, joint critics can be used in contrastive settings.
@@
Concatenated Quadratic Critic: We show that if $g$ and $h$ are linear projections (e.g., identity maps) and $f$ is a quadratic form of its concatenated arguments, then MI estimation, denoted here as $I_{\rm CCA}(X;Y)$, reduces to measuring canonical correlations (Hotelling, 1936) in a shared low-dimensional space (Appx. A.1.2).
Probabilistic Critic (VSIB Framework): Rather than using deterministic mappings, $f$, $g$, and $h$ can be stochastic, leading to variational objectives that regularize embedding distributions. We implement this with a loss similar to that introduced by Abdelaleem et al. (2025) (Appx. A.2):
@@ 4.1 Estimator Performance in the Infinite Data Regime
We first evaluate neural MI estimators in an idealized, effectively infinite-sample setting that eliminates overfitting. Although common in prior work, this regime obscures many real-world challenges. For benchmarking, we use both low- and high-dimensional synthetic data with known ground‑truth MI. We adopt the following notation. $X$ and $Y$ are the variables whose MI we estimate; $Z_{X}$ and $Z_{Y}$ denote their low‑dimensional embeddings by a critic. The observed dimensionalities are $K_{X}$ and $K_{Y}$ ($K$ if both are equal), while $K_{Z}$ is the true latent dimensionality when it differs from $K$. Finally, $k_{Z}$ is the embedding dimensionality of the critic, which may differ from $K_{Z}$.
@@
Figure 1: MI estimators in the low-dimensional, infinite-data regime. Each panel plots running MI estimates over training iterations for five true MI levels (increasing every 4000 iterations). Each step introduces a fresh batch of 128 samples. We compare the CCA-based estimator (optimal for Gaussian data), InfoNCE, SMILE, and their probabilistic variants (denoted with VSIB). Faint curves show raw estimates; bold curves show smoothed trends (Appx. A.6). Left: For jointly Gaussian $X,Y$, all estimators initially perform well. InfoNCE plateaus at its well-known intrinsic upper bound (van den Oord et al., 2018) $\log$(batch size) $\approx 7$ bits, while SMILE begins to overestimate at high MI, indicating overfitting. $I_{\rm CCA}$ overlaps with ground truth, as expected. Middle: Cubing $Y$ breaks linearity, and $I_{\rm CCA}$ fails. Nonetheless, InfoNCE behavior is almost unchanged, and SMILE remains reasonably effective with sufficient training (Appx. Fig. 10). Both slightly underestimate at low MI, and for SMILE this is largely offset by its intrinsic positive bias at high MI. Right: Passing $X$ and $Y$ through separate frozen teacher networks (one hidden softplus layer, 1024 units) creates highly nonlinear dependencies. All estimators underestimate MI.
@@
We begin with low-dimensional data ($K_{X}=K_{Y}=10$), where both traditional (e.g., kNN) and neural estimators are expected to perform well. We first analyze jointly Gaussian data with a chosen correlation. It is well known that, in this case, MI estimation reduces to the sum of information in each canonical correlation pair (Gelfand & Iaglom, 1959; Kullback, 1997; Huffmann & Mittelbach, 2022), $I_{\rm CCA}=-\frac{1}{2}\sum_{i}^{K_{Z}}\log(1-\rho_{i}^{2})$, where $\rho_{i}$ is the canonical correlation. We then increase complexity by (i) cubing $Y$ or (ii) passing both variables through separate fixed teacher networks, thereby introducing diverse nonlinear dependencies.
@@
Figure 1 (left) confirms that InfoNCE, SMILE, and CCA track the true MI on Gaussian data (until InfoNCE saturates (van den Oord et al., 2018)). The data matrices used here are aggregated across all batches seen during training at a given MI level. If the critic is expressive enough, SMILE is expected to show high variance and to overestimate at high MI since it evaluates log-sum-exp in its DV bound, Eq. (7), which is biased due to the Jensen’s inequality (Guo et al., 2022; Choi & Lee, 2020). We observe both problems, but neither is strong, at least at low MI.
@@
Cubing $Y$ introduces nonlinear correlations, and $I_{\rm CCA}$ fails, Fig. 1 (middle). In contrast, neural estimators recover MI after sufficient training (Appx. Fig. 10). Both exhibit a small but significant negative bias at low MI, likely due to the reduced invertibility of the cubic map near $y=0$. When $X$ and $Y$ are passed through separate frozen teacher networks, Fig. 1 (right), all estimators underestimate MI, although neural methods still outperform the CCA baseline. We again attribute this bias to the reduced invertibility, now due to the softplus saturation in the teacher networks.
@@
Appx. Fig. 8 surveys other neural estimators and confirms that InfoNCE and SMILE consistently outperform the rest. Therefore, in what follows, we restrict our analysis to these two methods, using a fixed clipping factor $\tau=5$ for SMILE. Combined, these results validate MI neural estimators—specifically InfoNCE, SMILE, and their VSIB variants—in low dimensions with abundant data. However, they still underestimate MI for strongly nonlinear dependencies, even with effectively unlimited data. Thus, seemingly, there is little reason to prefer them over simpler correlation‑based or kNN methods, which already perform well in these scenarios (Kraskov et al., 2004; Holmes & Nemenman, 2019; Czyz et al., 2023).
@@
Figure 2: MI estimators in the high-dimensional, infinite-data regime. We extend Fig. 1 to $K_{X}=K_{Y}=500$, embedding $K_{Z}=10$ latent variables into high-dimensional $X$ and $Y$. Left: A linear transformation (e.g., identity and replication) expands $Z$ to 500-dimensional $X$ and $Y$. $I_{\rm CCA}$ with $k_{z}\geq K_{Z}=10$ dimensions still accurately recovers the ground truth. Middle: A frozen nonlinear teacher network maps $Z$ to 500-dimensional $X$ and $Y$. Unlike the linear case, $I_{\rm CCA}$ fails due to the nonlinearity of the transformation. Increasing $k_{Z}>K_{Z}$ detects spurious correlations, inflating MI estimates and illustrating the limitations of linear methods in nonlinear settings. Right: Neural estimators (InfoNCE, SMILE, and their VSIB variants) are applied directly to the full 500-dimensional data. All are accurate across the full range of true MI values, performing even better than in Fig. 1 due to improved invertibility of the nonlinear transformation in high dimensions.
High-Dimensional $X$, $Y$.
@@
To distinguish effects of the observed and latent dimensionality, we increase the former, while fixing the latter. We start again with the jointly Gaussian 10-d $X$ and $Y$ with the known ground truth MI. We then (i) replicate each of the ten components of $X$ and $Y$ 50 times (denoted $X\to{\mathbf{I}}^{(50)\otimes}[X]$), and (ii) pass $X$, $Y$ through distinct frozen teacher networks, embedding each into 500 dimensions (denoted $X\to F_{X}^{(500)}[X]$, and similarly for $Y$). Both cases result in $K_{X}=K_{Y}\equiv K=500$, while $K_{Z}=10$, with the former case having only linear correlations.
@@
Figure 2 shows that, in the linear replication setup, $I_{\rm CCA}$ accurately recovers the true MI when $k_{Z}\geq K_{Z}$. In the nonlinear case, the CCA approach breaks down: as the number of detected canonical pairs $k_{Z}$ increases, the method approximates nonlinear dependencies with an ever larger set of linear projections, inflating the MI estimate with no upper bound. In contrast, neural estimators (InfoNCE, SMILE, and their VSIB variants) recover the ground truth MI when applied directly to the full $K=500$-dimensional spaces (up to InfoNCE’s saturation). Performance even surpasses the low-dimensional case, Fig. 1, because non-invertible, saturated softplus regions in one random embedding can be inverted in others, allowing reconstruction of the full latent manifold.
@@
This highlights two points. First, neural estimators may work well in high-dimensional nonlinear settings, where traditional approaches fail (Holmes & Nemenman, 2019) (Appx. Fig. 9). Second, matching the estimator to the structure of the data (e.g., using $I_{\rm CCA}$ for linear correlations) may result in more computationally and data-efficient estimation (Abdelaleem et al., 2024).